[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Jul 13, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Jul 14 03:12:39 UTC 2016


HSU Shift Log for Wed Jul 13 2016 23:12:33 GMT-0400 (EDT)

Shift: R

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: PAC-ATCF: when ATCF storm window was grabbed and panned/moved  
to the left or right sides of the monitor and the mouse button released,  
the window kept propagating on its own and nearly moved out of the monitor  
screen FOV. Rebooted PAC-ATCF, and issue appears to be resolved. Mouse  
function appears to quicker and more responsive, and panning of storm  
window stops after mouse is released.

N-AWIPS/AWIPS2 issues: None.

Other Techincal issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160714/f851bd4e/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list